Congratulations one more time. You have been appointed chair of Economic Advisers in Fantasyland. Income is currently $600,000, unemployment is 5 percent, and there are signs of coming inflation. You rely on your research assistant for specific numbers. He tells you that potential income is $564,000 and the mpe is 0.5.

a. The government wants to eliminate the inflationary gap by changing expenditures. What policy do you suggest?

(Click to select)IncreaseDecrease government spending by $ or (Click to select)increasedecrease taxes by $.

b. Your research assistant comes in and says “Sorry, I meant that the mpe is 0.8.” Redo your calculations for part a.

Instructions: Enter your responses rounded to the nearest whole dollar amount.

(Click to select)DecreaseIncrease government spending by $ or (Click to select)increasedecrease taxes by $.
Answer #1

a. When MPE is 0.5; the multiplier is 2 (=1/(1-0.5)). Because the economy is $36,000 above potential, thus advise to decrease the government spending by $18,000 (=36,000/ (1-0.5)) or increasing taxes by $36,000

b. When MPE is 0.8; the multiplier is 1.25 (=1/(1-0.2)). Because the economy is $36,000 above potential, thus advise to decrease the government spending by $7,200 (=36,000/ (1-0.8)) or increasing taxes by $9,000 (=7,200*1.25)